OpenStudy (anonymous):

Convert the polar equation r = sin(θ) sec2(θ) to Cartesian coordinates and describe the resulting curve.

OpenStudy (phi):

I assume you mean sec^2(θ) ? I would change the sec to 1/cos and multiply both sides to get r sin and r cos terms which are y and x in cartesian coords

OpenStudy (phi):

*mult by r

OpenStudy (phi):

can you try ?

OpenStudy (anonymous):

obviuosly its the one @phi suggested

OpenStudy (anonymous):

if it was not a square he was gonna put (2theta) and not 2(theta)

OpenStudy (phi):

Here is one step you could take \[ r = \sin(θ) \sec^2(θ) \] \[ r= \frac{r}{r} \frac{\sin(\theta)}{\cos(\theta)} \frac{1}{\cos(\theta) } \] you still have a cos to multiply by r, but if you multiply both sides by 1/r that will do it. now use x= r cos(A) y= r sin(A) to make it an equation in x and y